OpenStudy (anonymous):

What is the value of -3mn + 4m - 3 when m = 2 and n = -4? -43 -40 5 29

TheSmartOne (thesmartone):

Plug in m = 2 and n = -4 \(\sf\Large -3(2)(-4) + 4(2) - 3\)

TheSmartOne (thesmartone):

Remember, a negative number times a negative number equals a positive number

OpenStudy (anonymous):

i got -19 @TheSmartOne

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

-3(2)(-4) = -3(-4)(2) = 12(2) = 24

OpenStudy (anonymous):

24 + 8 -3 = 24 +5 = 29

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome.

TheSmartOne (thesmartone):

Thanks for helping while I was gone! @robtobey :)

TheSmartOne (thesmartone):

I was away from the computer for a while. So I was still online, but I wasn't actually here to help. :)

OpenStudy (anonymous):

Your welcome. BTY I did not know you were absent.

OpenStudy (anonymous):

Above posting corrected an earlier one of mine.